### v4

@@ 7.1 Adaptive and Automated Attacks
As LLMs become more powerful, attacks will also become more automated. Developing methods to systematically explore the space of possible prompt-jailbreaks or adversarial inputs (e.g. via generative tools) is an open challenge. Researchers must anticipate large-scale, automated exploit generation (self-playing AI attackers) and devise defenses that can scale accordingly [68].
@@ 7.2 Robust Alignment, Verification, and Control of Agentic LLMs
  • Scalable Oversight: Creating oversight mechanisms that can effectively monitor and intervene with highly autonomous and capable agents without stifling their utility.
@@ 7.5 Standardization and Collaboration
The community must establish security standards and best practices for LLM deployment. This includes benchmark suites for LLM robustness (especially against sophisticated agentic deception), shared threat models, and coordinated disclosure (e.g. companies and researchers sharing jailbreaks and novel deceptive behaviors so defenses can improve). As one recent survey suggests, developing efficient defense strategies and consensus guidelines is a priority for the field [17]. Collaboration between AI practitioners, security experts, and policymakers will be essential to keep pace with LLM advances.
@@ 7.6 Human-AI Interaction Research
LLMs interact with users in novel ways. Studying how humans can detect or guard against malicious AI outputs, designing user interfaces that highlight AI uncertainties or potential deceptiveness, and ensuring accountability are open areas.
@@ 7.7 Emerging Proactive Measures and Ongoing Efforts
Progress in these directions is already underway, with new multi-layer safeguards emerging. For example, Meta’s Llama Guard 3 combines a policy LLM and a vision encoder to filter both text and images before they reach the main model, achieving 99.4% precision on the Harassment/Hate category [12]. Similarly, chain-of-utterances red-teaming automates the discovery of multi-turn failure cases and has already uncovered jailbreaks missed by single-turn probes [9]. Nonetheless, as evaluations across efforts such as BackdoorLLM and RAG safety studies confirm, current defenses often remain piecemeal, and attackers continue to adapt quickly, underscoring the ongoing nature of these challenges.
@@ 8 Conclusion
The advent of LLMs brings both unprecedented AI capabilities and new security risks. This survey has outlined the main threat categories – from prompt hijacking and adversarial manipulation to malicious use cases and the profound challenges posed by autonomous agent hazards such as goal misalignment, emergent deception, scheming, and self-preservation. We have shown that while a variety of defenses have been proposed, they currently offer only partial protection, and may be ineffective against more sophisticated, internally motivated deceptive behaviors that can persist through current safety training. As AI systems grow more capable and autonomous, security concerns will not only persist but are also likely to intensify, becoming a long-term challenge that evolves in tandem with AI progress. The open challenges ahead are daunting but clear: we must develop more effective defenses, rigorous alignment and verification methods capable of addressing strategic agentic deception, and industry-wide standards for LLM security. As LLMs continue to proliferate in critical applications, it is crucial for the AI and security communities to prioritize safety and control. By understanding and mitigating these risks preemptively, especially those related to the potential for autonomous LLM agents to develop and pursue their own covert intentions, we can help ensure that powerful LLM technology remains safe, secure, and beneficial for society.

### v5

@@ 7.1 Adaptive and Automated Attacks
As LLMs become more powerful, attacks will also become more automated. Developing methods to systematically explore the space of possible prompt injections is an open challenge. Researchers must anticipate large-scale, automated exploit generation (self-playing AI attackers) and devise defenses that can scale accordingly [70].
@@ 7.2 Robust Alignment, Verification, and Control of Agentic LLMs
  • Scalable Oversight: Creating oversight mechanisms that can effectively monitor and intervene with highly autonomous and capable agents without stifling their utility. ne promising, though still theoretical, direction is the development of non-agentic paradigms such as the proposed ’Scientist AI’ [8], which would serve as a trustworthy oversight system by design rather than through reactive intervention.
@@ 7.5 Standardization and Collaboration
The community must establish security standards and best practices for LLM deployment. This includes benchmark suites for LLM robustness (especially against sophisticated agentic deception), shared threat models, and coordinated disclosure (e.g. companies and researchers sharing jailbreaks and novel deceptive behaviors so defenses can improve). As a survey from 2024 suggests, developing efficient defense strategies and consensus guidelines is a priority for the field [17]. Collaboration between AI practitioners, security experts, and policymakers will be essential to keep pace with LLM advances.
@@ 7.6 Human-AI Interaction Research
LLMs interact with users in novel ways as their capabilities evolve. Studying how humans can detect or guard against malicious AI outputs, designing user interfaces that highlight AI uncertainties or potential deceptiveness, and ensuring accountability are open areas.
@@ 7.7 Emerging Proactive Measures and Ongoing Efforts
Progress in these directions is already underway, with new multi-layer safeguards emerging. For example, Meta’s Llama Guard 3 combines a policy LLM and a vision encoder to filter both text and images before they reach the main model, achieving 99.4% precision on the Harassment/Hate category [12]. Similarly, chain-of-utterances red-teaming automates the discovery of multi-turn failure cases and has already uncovered jailbreaks missed by single-turn probes [9]. Nonetheless, as evaluations across efforts such as BackdoorLLM and RAG safety studies confirm, current defenses often remain piecemeal, and attackers continue to adapt quickly, underscoring the ongoing nature of these challenges [35, 47].
@@ 8 Conclusion
The advent of LLMs brings both unprecedented AI capabilities and new security risks. This survey has outlined the main threat categories – from inference-time and training-time attacks to malicious use cases and the profound challenges posed by autonomous agent hazards. We have shown that while a variety of defenses have been proposed, each of them currently offers only partial protection, and may be ineffective against more sophisticated, internally motivated deceptive behaviors that can persist through current safety training. As AI systems grow more capable and autonomous, security concerns will not only persist but are also likely to intensify, becoming a long-term challenge that evolves in tandem with AI progress. The open challenges ahead are daunting but clear: we must develop more effective defenses, rigorous alignment and verification methods capable of addressing strategic agentic deception, and industry-wide standards for LLM security. As LLMs continue to proliferate in critical applications, it is crucial for the AI and security communities to prioritize safety and control. By understanding and mitigating these risks preemptively, especially those related to the potential for autonomous LLM agents to develop and pursue their own covert intentions, we can help ensure that powerful LLM technology remains safe, secure, and beneficial for society.